OpenStudy (anonymous):

carlos took a cab home from the airport. the total cost of the cab ride was $19.00. write a formula that helps carlos determine how much he paid per mile. lets let P=price per mile n=number of miles traveled explain.

TheSmartOne (thesmartone):

pn=19 I am not sure..

OpenStudy (swissgirl):

P*n=$19.00

TheSmartOne (thesmartone):

I think it is pn=19

TheSmartOne (thesmartone):

Yup @swissgirl

OpenStudy (swissgirl):

Most cabs dont charge like this though -.-

OpenStudy (swissgirl):

They charge you a starting fee and then to add to that that charge you a fee per mile

OpenStudy (anonymous):

well its a question lol idk

OpenStudy (swissgirl):

hahaha im just teasing

OpenStudy (anonymous):

oh okay i was starting to get lost like i was wondering if i did put the problem wrong

TheSmartOne (thesmartone):

lol

OpenStudy (anonymous):

well i still dont know how to solve it :p

TheSmartOne (thesmartone):

Well if you travel p miles then n miles is how much each mile costs.

OpenStudy (anonymous):

still lost v.v

OpenStudy (swissgirl):

ok lemme show you

OpenStudy (swissgirl):

SO basically lemme show you this with an example So for every mile the cab drives he charges you 30 cents. Now in total you pay 19 dollars so 0 .30*Amount of miles drive =19.00

OpenStudy (swissgirl):

so let n= the amount of miles you drove .30*n=19

OpenStudy (anonymous):

oh so 19.00 *0.30?

OpenStudy (swissgirl):

nope nope n refers to the amount of miles he drove

OpenStudy (swissgirl):

19 is the amount he paid for the taxi

OpenStudy (anonymous):

yeah so i either divide or mutiply by 0.30 right???

OpenStudy (swissgirl):

.30*n=19 Divide .30 from both sides n=19/.30

OpenStudy (swissgirl):

Ok here we used that p=.30 But in your original problem we didnt specify how much p is we just said that p represent the amount of money the cab would charge per mile so that is why its P*n=19

OpenStudy (anonymous):

so then 19/0.30?

OpenStudy (swissgirl):

yup But made up that p=.30 Your problem didnt specify what p is

OpenStudy (anonymous):

so its 63?

OpenStudy (swissgirl):

ya approximately 63.33 But this isnt what ur question asked for I just gave you an example so that you can understand the concept This aint the answer to your question

OpenStudy (anonymous):

uhhh i hate word questions my head hurts

OpenStudy (swissgirl):

Your question asked you for the formula THATS IT So 19=p*n

OpenStudy (anonymous):

19 * .30????

OpenStudy (swissgirl):

ahhhh its gonna be P*n=19 19/n=p so the final answer is 19/n=p